# A Survey on Vision-Language-Action Models for Embodied AI

### v5

@@ III-B9 Large VLA
RT-2 [2] endeavors to harness the capabilities of large multimodal models in robotics tasks, drawing inspiration from models like PaLI-X and PaLM-E. The approach introduces co-fine-tuning, aiming to fit the model to both Internet-scale visual question answering (VQA) data and robot data. This training scheme enhances the model’s generalizability and brings about emergent capabilities.
@@
RT-X [112] builds upon the previous RT-1 and RT-2 models. These models are re-trained using the newly introduced open-source large dataset, named Open X-Embodiment (OXE), which is orders of magnitude larger than previous datasets. The resulting models, RT-1-X and RT-2-X, both outperform their original versions.
@@
$\pi_{0}$ [115] proposes a flow-matching architecture for transforming VLMs into VLAs. By incorporating an additional action expert based on the mixture-of-experts framework, it effectively inherits the Internet-scale knowledge in the VLM while extending its capabilities to address robotic tasks.
@@
Visual autoregressive modeling (VAR) [133] with quantized visual tokens demonstrates improved performance over diffusion models in image generation. This suggests that the three modalities of VLAs can be unified under the autoregressive paradigm [134]. WorldVLA [122] and UniVLA [123] advance this direction by integrating VLAs with world models. It quantizes multimodal data into discrete tokens, forming a shared vocabulary of quantized multimodal tokens. Consequently, all modalities can be modeled autoregressively, enabling not only action and text generation but also image generation, thereby constituting a world model.
@@
Representative VLA architectures are illustrated in Figure 5. FiLM is used in RT-1 and thus its follow-up models inherit this mechanism. While cross-attention may offer superior performance with smaller model sizes, concatenation is simpler to implement and can achieve comparable results with larger models [126]. Quantization unifies multimodal tokens into a shared vocabulary, thus enabling integration with world models. The tool-use paradigm of LLMs can also be applied to robotic tasks.
@@
The behavior cloning (BC) objective is used in imitation-learning, with different variants for different action types. The BC objective for continuous action [99] can be written as:
@@
Discrete action is achieved by dividing the action value range into a number of bins [94]. Its BC objective is:
@@
CLIPort [31] and VIMA [126] use $SE(2)$ action and its BC objective can be expressed as follows:
@@
The DDPM objective in diffusion-based control policies [104] is represented as:
@@
While discrete action has demonstrated superior performance in RT-1 [94], Octo [107] argues that it leads to early grasping issues. $SE(2)$ actions require the model to predict only the pick and place end-effector poses, which is sufficient for many tabletop manipulation tasks. However, more complex tasks—such as “pouring water into a cup”—may require additional DoFs, thereby necessitating $SE(3)$ actions [38]. Although point-based actions can be coarse-grained, they are more easily obtained from VLMs in a zero-shot manner [38].
@@
While LVLAs can greatly enhance instruction-following abilities because they can better parse user intentions, concerns arise regarding their training cost and deployment speed. Slow inference speed, in particular, can significantly impact performance in dynamic environments, as changes in the environment may occur during inference. Therefore, several methods have been proposed to improve efficiency. TinyVLA [118] focused on inference speed and data efficiency through smaller VLM and diffusion head for robot action. DeeR-VLA [101] proposes only partially activating the model with dynamic inference with early-exit.
@@
Similar to LLMs, scaling laws have also been observed in robotics [149], revealing the importance of model size, data quality, and diversity in environments and objects. Further research in this area can guide the development of VLAs with improved in-the-wild generalization capabilities.
@@ IV-A2 End-to-end Task Planners with 3D Vision
Some task planners also explore the use of 3D vision. Because the majority of current MLLMs deal with images as visual inputs, they require architectural changes to incorporate 3D visual inputs, and therefore they are usually end-to-end models.
@@ IV-B1 Language-based Task Planners
Language-based approaches use natural language descriptions as the medium for exchanging multimodal information, as illustrated in Figure 6a.
Inner Monologue [9] sits between high-level command and low-level policy to enable closed-loop control planning. It employs LLM to generate language instructions for low-level control policies and dynamically updates these instructions based on feedback received from the control policies. The feedback encompasses various sources: success feedback, object and scene feedback, and human feedback. As the feedback is communicated to the LLM in textual format, no additional training is required for the LLM. A similar approach is used in ReAct [76].
@@
Socratic Models (SMs) [143] presents a unique framework wherein diverse pretrained models are effectively composed without the need for finetuning. The framework is based on the key component, named multimodal-informed prompting, facilitating information exchange among models with varied multimodal capabilities. The idea is to utilize multimodal models to convert non-language inputs into language descriptions, effectively unifying different modalities within the language space. Beyond excelling in conventional multimodal tasks, SMs showcases its versatility in robot perception and planning. In addition to natural language plans, task plans can also be represented in the form of pseudocode.
@@ IV-B2 Code-based Task Planners
ChatGPT for Robotics [145] take advantage of the programming ability of ChatGPT to facilitate “user on the loop” control, a departure from the conventional “engineer in the loop” methodology. The procedure includes several steps: firstly, a list of APIs is defined, such as an object-detection API, a grasp API, a move API; secondly, a prompt is then constructed for ChatGPT, specifying the environment, API functionality, task goal, etc.; thirdly, iteratively prompting ChatGPT to write code with the defined APIs that can execute the task, provided the access to simulation and user feedback for evaluating the code quality and safety; finally, executing the ChatGPT generated code. In this procedure, ChatGPT serves as a high-level task planner and actions are generated through function calls to corresponding low-level APIs.
Code as policies (CaP) [146] also leverages the code-writing capability of LLMs. It employs GPT-3 or Codex to generate policy code, which, in turn, invokes perception modules and control APIs. CaP exhibits proficiency on spatial-geometric reasoning, generalization to new instructions, and parameterization for low-level control primitives. By leveraging the multimodal capabilities of GPT-4V, COME-robot [152] eliminates the need for perception APIs in CaP. This also opens up possibilities for open-ended reasoning and adaptive planning within a closed-loop framework, enabling capabilities such as failure recovery and free-form instruction following.
DEPS [147] stands for “Describe, Explain, Plan and Select”. This approach employs an LLM to generate plans and explain failures based on feedback descriptions collected from the environment—a process referred to as “self-explanation”, aiding in re-planning. Additionally, DEPS introduces a trainable goal selector to choose among parallel candidate sub-goals based on how easily they can be achieved, a crucial aspect often overlooked by other high-level task planners.
@@ V Datasets and Benchmarks
V-A Real-world Robot Datasets & Benchmarks
V-B Simulators, Simulated Robot Datasets & Benchmarks
@@
VirtualHome [173] 7/- -/509 Lang Unity Navi, Mani RGB, D, S Force, PD Human Object States, Task Planning LID, Translated〈LM〉, ProgPrompt
@@ V-D Human Datasets
An alternative strategy to address data scarcity in real-world settings is to leverage human data. Human behavior offers plentiful guidance for robot policies due to its dexterity and diversity. However, this strategy also comes with inherent drawbacks. Capturing and transferring human hand/body motions to robot embodiments is inherently difficult. Moreover, the inconsistency in human data poses a hurdle, as some data may be egocentric while others are captured from third-person perspectives. Additionally, filtering human data to extract useful information can be labor-intensive. These obstacles underscore the complexities involved in incorporating human data into robot learning processes. UMI [190] proposes a method to mitigate these issues using hand-held grippers. For a more comprehensive comparison of human datasets, we refer interested readers to [191].
@@ V-F Embodied Question Answering Benchmarks
Embodied question answering (EQA) benchmarks, as summarized in Table VII, do not directly evaluate robotic tasks like manipulation and navigation, but they are aimed at other relevant abilities for embodied AI, such as spatial reasoning, physics understanding, as well as world knowledge. EQA is akin to previous visual question answering benchmarks, but differs in that the agent can actively explore the environment before providing an answer. EmbodiedQA [197] and IQUAD [198] were among the first works to introduce this type of benchmarks. MT-EQA [199] focuses on complex questions involving multiple targets. MP3D-EQA [200] converts previous RGB inputs to point clouds, testing 3D perception capabilities.
@@ VI Challenges and Future Directions
VLA foundation models or robotic foundation models (RFM) for embodied AI remains an open research topic primarily due to the diversity in embodiments, environments, and tasks. Many have made significant progress [110, 115], but still lacks generalization capability on par with LLMs in NLP. Attaining such a level of generalization is very challenging, as it requires the development of many core AGI capabilities.

### v7

@@ III-B9 Large VLA
RT-2 [1] endeavors to harness the capabilities of large multimodal models in robotics tasks, drawing inspiration from models like PaLI-X and PaLM-E (see the architecture in Fig. 5e). The approach introduces co-fine-tuning, aiming to fit the model to both Internet-scale visual question answering (VQA) data and robot data. This training scheme enhances generalizability and brings about emergent capabilities.
@@
RT-X [113] builds upon the previous RT-1 and RT-2 models. These models are retrained using the newly introduced open-source large dataset, named Open X-Embodiment (OXE), which is orders of magnitude larger than previous datasets. The resulting models, RT-1-X and RT-2-X, both outperform their original versions.
@@
$\pi_{0}$ [116] proposes a flow-matching architecture for transforming VLMs into VLAs. By incorporating an additional action expert based on the mixture-of-experts framework (see the architecture in Fig. 5f), it effectively inherits the Internet-scale knowledge in the VLM while extending its capabilities to address robotic tasks.
@@
GR00T N1 [133] introduces a dual-system architecture to build a robot foundation model for humanoid robots (see the architecture in Fig. 5g). Its VLM-based System 2 processes image observations and language instructions at 10 Hz, while System 1 is a diffusion module that generates closed-loop motor actions in real time (120 Hz).
NORA-1.5 [134] unifies a VLA with a world model through reward-guided post-training. Genie Envisioner [135] is a world foundation platform that integrates a world model and a VLA within a single video-generative framework. See the general architecture in Fig. 5h.
Visual autoregressive modeling (VAR) [136] with quantized visual tokens demonstrates improved performance over diffusion models in image generation. This suggests that the three modalities of VLAs can be unified under the autoregressive paradigm [137]. WorldVLA [124] and UniVLA [125] advance this direction by integrating VLAs with world models (see the architecture in Fig. 5i). It quantizes multimodal data into discrete tokens, forming a shared vocabulary of quantized multimodal tokens. Consequently, all modalities can be modeled autoregressively, enabling not only action and text generation but also image generation, thereby constituting a world model.
@@
Representative VLA architectures are illustrated in Figure 5. FiLM is used in RT-1, and thus its follow-up models inherit this mechanism. While cross-attention may offer superior performance with smaller model sizes, concatenation is simpler to implement and can achieve comparable results with larger models [128]. Quantization unifies multimodal tokens into a shared vocabulary, thus enabling integration with world models. The tool-use paradigm of LLMs can also be applied to robotic tasks.
@@
The behavior cloning (BC) objective is used in imitation-learning, with different variants for different action types. The BC objective for continuous action can be written as:
@@
Discrete action is achieved by dividing the action value range into a fixed number of bins. Its BC objective is:
@@
The BC objective of $SE(2)$ action is:
@@
The DDPM objective in diffusion-based control policies is:
@@
While discrete action has demonstrated superior performance in RT-1 [95], Octo [108] argues that it leads to early grasping issues. $SE(2)$ actions require the model to predict only the pick and place end-effector poses, which is sufficient for many tabletop manipulation tasks. However, more complex tasks—such as “pouring water into a cup”—may require additional DoFs, thereby necessitating $SE(3)$ actions [35]. Although point-based actions can be coarse-grained, they are more easily obtained from VLMs in a zero-shot manner.
@@
While LVLAs can greatly enhance instruction-following abilities because they can better parse user intentions, concerns arise regarding their training cost and deployment speed. Slow inference speed, in particular, can significantly impact performance in dynamic environments, as changes in the environment may occur during inference. Therefore, several methods have been proposed to improve efficiency. TinyVLA [119] focused on inference speed and data efficiency through a smaller VLM and diffusion head for robot action. DeeR-VLA [102] proposes only partially activating the model with dynamic inference with early-exit.
@@
Similar to LLMs, scaling laws have also been observed in robotics [151, 152], revealing the importance of model size, dataset size, and the diversity of environments and objects. Further research in this area can guide the development of VLAs with robust in-the-wild generalization capabilities.
@@ IV-A2 End-to-end Task Planners with 3D Vision
Some task planners also explore the use of 3D vision. Because the majority of current MLLMs deal with images as visual inputs, they require architectural changes to incorporate 3D visual inputs, and therefore, they are usually end-to-end models.
@@ IV-B1 Language-based Task Planners
Language-based approaches use natural language descriptions as the medium for exchanging multimodal information, as shown in Figure 6a.
Inner Monologue [8] sits between high-level command and low-level policy to enable closed-loop control planning. It employs LLM to generate language instructions for low-level policies and dynamically updates these instructions based on feedback received from them. The feedback encompasses various sources: success feedback, object and scene feedback, and human feedback. As the feedback is communicated to the LLM in textual format, no additional training is required for the LLM. A similar approach is used in ReAct [74].
@@
Socratic Models (SMs) [145] present a unique framework wherein diverse pretrained models are effectively composed without the need for finetuning. The framework is based on the key component, named multimodal-informed prompting, facilitating information exchange among models with varied multimodal capabilities. The idea is to utilize multimodal models to convert non-language inputs into language descriptions, effectively unifying different modalities within the language space. Beyond excelling in conventional multimodal tasks, SMs showcase their versatility in robot perception and planning. In addition to natural language plans, task plans can also be represented in the form of pseudocode.
@@ IV-B2 Code-based Task Planners
ChatGPT for Robotics [147] takes advantage of the programming ability of ChatGPT to facilitate “user on the loop” control, a departure from the conventional “engineer in the loop” methodology. The procedure includes several steps: firstly, a list of APIs is defined, such as an object-detection API, a grasp API, a move API; secondly, a prompt is then constructed for ChatGPT, specifying the environment, API functionality, task goal, etc.; thirdly, iteratively prompting ChatGPT to write code with the defined APIs that can execute the task, provided the access to simulation and user feedback for evaluating the code quality and safety; finally, executing the ChatGPT generated code. In this procedure, ChatGPT serves as a high-level task planner, and actions are generated through function calls to corresponding low-level APIs.
Code as policies (CaP) [148] also leverages the code-writing capability of LLMs. It employs GPT-3 or Codex to generate policy code, which, in turn, invokes perception modules and control APIs. CaP exhibits proficiency in spatial-geometric reasoning, generalization to new instructions, and parameterization for low-level control primitives. By leveraging the multimodal capabilities of GPT-4V, COME-robot [155] eliminates the need for perception APIs in CaP. This also opens up possibilities for open-ended reasoning and adaptive planning within a closed-loop framework, enabling capabilities such as failure recovery and free-form instruction following.
DEPS [149] stands for “Describe, Explain, Plan, and Select”. This approach employs an LLM to generate plans and explain failures based on feedback descriptions collected from the environment—a process referred to as “self-explanation”, aiding in re-planning. Additionally, DEPS introduces a trainable goal selector to choose among parallel candidate sub-goals based on how easily they can be achieved, a crucial aspect often overlooked by other high-level task planners.
@@ V Datasets and Benchmarks
V-A Real-world Robot Datasets & Benchmarks
V-B Simulators, Simulated Robot Datasets & Benchmarks
@@
VirtualHome [176] 7/- -/509 Lang Unity Navi, Mani RGB, D, S Force, PD Human Object States, Task Planning LID, Translated⟨LM⟩, ProgPrompt
@@ V-D Human Datasets
An alternative strategy to address data scarcity in real-world settings is to leverage human data. Human behavior offers plentiful guidance for robot policies due to its dexterity and diversity [193]. However, this strategy also comes with inherent drawbacks. Capturing and transferring human hand/body motions to robot embodiments is inherently difficult. Moreover, the inconsistency in human data poses a hurdle, as some data may be egocentric while others are captured from third-person perspectives. Additionally, filtering human data to extract useful information can be labor-intensive. These obstacles underscore the complexities involved in incorporating human data into robot learning processes. UMI [194] proposes a method to mitigate these issues using hand-held grippers. For a more comprehensive comparison of human datasets, we refer interested readers to [195].
@@ V-F Embodied Question Answering Benchmarks
Embodied question answering (EQA) benchmarks, as summarized in Table VII, do not directly evaluate robotic tasks like manipulation and navigation, but they are aimed at other relevant abilities for embodied AI, such as spatial reasoning, physics understanding, as well as world knowledge. EQA is akin to previous visual question answering benchmarks, but differs in that the agent can actively explore the environment before providing an answer. EmbodiedQA [201] and IQUAD [202] were among the first works to introduce this type of benchmark. MT-EQA [203] focuses on complex questions involving multiple targets. MP3D-EQA [204] converts previous RGB inputs to point clouds, testing 3D perception capabilities.
@@ VI Challenges and Future Directions
VLA foundation models or robotic foundation models (RFM) for embodied AI remain an open research topic primarily due to the diversity in embodiments, environments, and tasks. Many have made significant progress [111, 116], but still lack generalization capability on par with LLMs in NLP. Attaining such a level of generalization is very challenging, as it requires the development of many core AGI capabilities.
@@ VII Conclusion
Acknowledgments
The research presented in this paper was partially supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (CUHK 2410072, RGC R1015-23) and CUHK 7010870.